OpenStudy (anonymous):

a cat weighs 10 pounds plus half its weight. how much does the cat weigh? a cat weighs 10 pounds plus half its weight. how much does the cat weigh? @Mathematics

OpenStudy (anonymous):

10+1/2w

OpenStudy (anonymous):

15 pounds

OpenStudy (anonymous):

10+1/2(10)=15

OpenStudy (anonymous):

I think its 20 pounds w=10+1/2w 1/2w=10 w=20

OpenStudy (anonymous):

The question is worded in a way that there are a lot of interpretations so your guesses are as good as mine.

OpenStudy (anonymous):

a cat weighs 10 pounds (10) plus half its weight +1/2(10) . how much does the cat weigh? =w put it together now. 10+1/2(10)=w w=15 Andras, the equation you have created will end up having w^2 because you will move it to the other side. you will end up w/ an irrational # :) I think lol

OpenStudy (anonymous):

I still stick with my solution. And it wont have any w^2

OpenStudy (anonymous):

15 pounds

OpenStudy (anonymous):

No Andras is right... think about the question and see if your answer makes sense... 10 pounds plus half its weight.. if it is 15 like you say then its 10 + 7.5, but that doesn't equal 15... but with 20: 10 + half 20 = 20... so it does makes sense

OpenStudy (anonymous):

i checked, you are right. sorry i forgot that you'll subtract. but i still say 15 just bcuz of the way its worded

OpenStudy (anonymous):

half of 10 is 5 anyway

OpenStudy (anonymous):

yeah, how did 7.5 come out of these #s?

OpenStudy (anonymous):

I have no clue

OpenStudy (anonymous):

15/2 you said 15 was the weight

OpenStudy (anonymous):

but thats half of 15

OpenStudy (anonymous):

yes half the weight like the question says

OpenStudy (anonymous):

so 15 is wrong

OpenStudy (anonymous):

it's asking half the weight of 10lbs.

OpenStudy (anonymous):

it's saying PLUS

OpenStudy (anonymous):

5lbs.

OpenStudy (anonymous):

therefor the answer is 15lbs.

OpenStudy (anonymous):

no its asking for half the weight not half of 10..

OpenStudy (anonymous):

thank you!! lol

OpenStudy (anonymous):

what if Pinnoccio said " And now my nose will grow"?

OpenStudy (anonymous):

What?